40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 21     Entry time: Sat Oct 27 19:00:44 2007
Author: waldman 
Type: Configuration 
Category: OMC 
Subject: Hanging, locked OMC with REFL extracted. 
I got the OMC locked to the fiber output today. It was much more difficult than I expected and I spent about 30 minutes or so flailing before stopping to think. The basic problem is that the initial alignment is a search in 4-dimensional space and there is naturally only one signal, the reflected DC level, to guide the alignment. I tried to eyeball the alignment using the IR card and "centering" the beams on mirrors, but I couldn't get close enough to get any light through. I also tried to put a camera on the high reflector transmission, but with 1.5 mW incident on the cavity, there is only 1.5 microwatts leaking through in the best case scenario, and much, much less during alignment.

I resolved the problem by placing a high reflector on a 3.5 inch tall fixed mount and picking off the OMC transmitted beam before it reaches the DC diodes. I took the pickoff beam to a camera. The alignment still sucked because even though the beam cleanly transmitted the output coupler, it wasn't anywhere close to getting through the OTAS. To resolve this problem, I visually looked through the back of M2 at M1 and used the IR card to align the beam to the centers of each mirror. That was close enough to get me fringes and align the camera. With the camera aligned, the rest was very easy.

I restored the PDH setup we know and love from the construction days and locked the laser to the OMC with no difficulty. The laser is in Rana's lab so I send the +/- 10V control signal from the SR560 down a cable to 058E where it goes into the Battery+resistor box, the Throlabs HV amplifier, and finally the FAST channel of the NPRO. BTW, a simple experiment sows that about 35 +/- 3 V are required to get an FSR out of the NPRO, hence the Thorlabs HV. The EOM, mixer, splitter, etc is on the edge of the table.

With this specific OMC alignment, ie. the particular sitting on EQ stops, it looks like all of the ghost beams have a good chance of coming clear. I can fit a 2 inch optic in a fixed mount in between the end of the breadboard and the leg of the support structure. A picture might or might not be included someday. One of the ghost beams craters directly into the EQ stop vertical member. The other ghost barely misses M2 on its way down the length of the board. In its current configuration, the many REFL beam misses the leg by about 1.5 inches.
ELOG V3.1.3-